# PSBD: Prediction Shift Uncertainty Unlocks Backdoor Detection

### v1

@@ Appendix A Pilot studies on More Attacks
In this section, we will show the more results of our pilot study in Section 4.1 and a combined method between MC-Dropout and SCP. We maintain the same experimental setup with pilot study that conduct experiments on the CIFAR-10 dataset with ResNet-18 trained 100 epochs. Apart from Adaptive-Blend attack, which has a poisoning ratio of 1%, all other attacks maintain a poisoning ratio of 10%.
@@
(a) Blend
(b) TrojanNN
(c) Label-Consistence
(d) ISSBA
(e) Adaptive-Blend
Figure A1: The average MC-Dropout uncertainty of clean training data, backdoor training data and clean validation data under various poisoned models.
@@
In the Section  4.1 of our main paper, we can observe that the model’s mapping from trigger to target label in backdoor data is more salient and robust compared to general image features. In light of this observation, we hypothesize the potential factor of the proximity between backdoor training data uncertainty and clean training data under WaNet attack can be attributed to the absence of sufficient sabotage on the image features. We expect that the model’s ability to extract feature patterns from clean data will be significantly affected, while that from backdoor data will be minimally impacted. Consequently, we can enlarge the uncertainty gap between clean data and backdoor data. One direct method is to increase the dropout rate $p$. However, due to our inability to ascertain the presence or quantity of backdoor data within the suspicious training dataset, we encounter difficulty in determining an optimal value for $p$. Thus, we incorporate elements of SCP approach to introduce a more controllable uncertainty with MC-Dropout uncertainty. SCP can be considered as an instance of input-level uncertainty. It amplifies the pixel values of input images by multiple times, aiming to disrupt the feature patterns in the image. Therefore, we conduct further experiments to see if adding input-level uncertainty further disrupts the general feature patterns in the image.
@@ Results.
However, under WaNet scenario, one can see from Figure 2(f) that although the uncertainty of the clean training data becomes closer to that of the clean validation data after scaling, the uncertainty gap between clean training and backdoor training data further reduces. Furthermore, as shown in Figure 2(e) and Figure 2(h), under the Label-Consistence and the Adaptive-Blend attack scenarios, the backdoor training data exhibits only slightly smaller uncertainty than that of clean training data, which poses significant challenges for their differentiation.
Appendix B Prediction Shift Phenomenon on More Scenarios
B.1 PS Phenomenon on More Poisoned Models.
@@ Appendix B Prediction Shift Phenomenon on More Scenarios
Except for the ISSBA attack, the shift intensity of clean data is not pronounced. In other attacks, the shift intensity of clean data is extremely strong. More importantly, in all attack scenarios, clean data exhibits a bias towards the target class (class 0 in our experiments). This further indicates that when the model has good generalizability, the neuron bias path established by the backdoor data in the model becomes more stable and specific. This property may be exploited in the future to detect the target class of backdoor attacks.
B.2 PS Phenomenon on Tiny ImageNet Dataset.
@@
From the observation of Figure A4, we can find that the PS phenomenon and neuron bias effect persist even in the more complex Tiny ImageNet dataset.
@@
(a) Benign Model
(b) BadNets
(c) WaNet
Despite assuming that the defender can freely choose the model architecture, we also conducted experiments using the VGG16-bn model to show that our method is not dependent on any specific model architecture. The experiment was conducted as follows Section 4.2.
@@
Appendix C Experiments Details
C.1 Detailed Settings for Datasets and Training of Backdoored Models.
@@ Appendix C Experiments Details
Dataset # Input size Classes Training images Testing images Validation images
CIFAR-10 3$\times$32$\times$32 10 50,000 10,000 2,500
Tiny ImageNet 3$\times$64$\times$64 200 100,000 10,000 5,000
@@
The details of datasets and training procedures of DNN models in our experiments are summarized in Table A1 and Table A2. Note that clean validation data refers to the clean data we used to filter backdoor training data, which was 5% of the total quantity of whole training dataset randomly selected from the test set of CIFAR-10 and the validation set of Tiny ImageNet, respectively.
C.2 Implements of Backdoor Attacks.
BadNets, TrojanNN, and Blend correspond to typical all-to-one label-poisoned attacks with patch-like trigger, generated trigger, and blending trigger respectively. Label-Consistent is a representative clean label attack. WaNet is an image transformation-based invisible attack. ISSBA is an effective sample-specific invisible attack which generates sample-specific invisible additive noises as backdoor triggers. It generates sample-specific invisible additive noises as backdoor triggers by encoding an attacker-specified string into benign images through an encoder-decoder network. Adaptive-Blend is an adaptive poisoning strategy suggested that can suppress the latent separability characteristic.
@@
We implemented this attack using the backdoor-toolbox. The trigger we used on CIFAR-10 is a 3$\times$3 checkerboard placed in the bottom right corner of an image, and a 6$\times$6 trigger placed in the same position on Tiny ImageNet.
@@
On CIFAR-10, we directly use the adversarial images provided by the original paper; [^3] https://github.com/MadryLab/label-consistent-backdoor-code on Tiny ImageNet we use the adversarial images provided by BackdoorBench. It should be noted that the data provided by BackdoorBench does not have triggers added to the corners of the images, as described in the original paper. Therefore, we have added triggers to the original data to maintain consistency with the original paper [24]. The transparency of the triggers was set to 0.05.
@@
We implemented this attack using the backdoor-toolbox. Following the original paper [38], we choose “Hello Kitty” trigger and set the the cover ratio equal to the poisoning ratio. Compared to the original paper [38], we only added the cover ratio and poisoning ratio to ensure that the attack success rate exceeds 85%. On CIFAR-10, we selected the “Hello Kitty” trigger, setting both the cover ratio and poisoning ratio to 0.01, and the blend ratio to 0.2. On Tiny ImageNet, we chose a random noise trigger, setting both the cover ratio and poisoning ratio to 0.02, and the blend ratio to 0.15.
C.3 Performance of the Benign and Poisoned Models.
Consistent with the methodology employed in previous backdoor attack studies, we utilize performance metrics to assess the effectiveness of the backdoor attacks: attack success rate (ASR) and clean accuracy (CA). ASR denotes the success rate in classifying the poisoned samples into the corresponding target classes. CA measures the accuracy of the backdoored model on the benign test dataset. ASR and CA for different backdoor attacks are included in Table A3.
CA ASR   CA ASR   CA ASR   CA ASR   CA ASR   CA ASR   CA ASR   CA
C.4 Implements of Baseline Defences.
We implement Spectral Signature [39], Strip[19], Spectre[40] and SCAN [41] based on the original implementation provided by the backdoor-toolbox. We have implemented the SCP [18] based on the backdoor-toolbox.
C.5 Computational Environment
@@
Appendix D Detection Performance on Low Poisoning Ratio Scenario
@@
As shown in Table A4, PSBD demonstrated superior performance against various backdoor attacks under low poisoning scenarios, outperforming all other baseline methods on average. However, its performance exhibited a slight decline in certain attack settings. This degradation may be attributed to the reduced robustness of neuron bias paths within the model due to the limited amount of backdoor data, making it more susceptible to random fluctuations. The performance of other baseline methods deteriorated significantly, particularly on the more challenging Tiny ImageNet dataset. While Spectre achieved high TPR on CIFAR-10, its excessive FPR renders it impractical. SCAN, despite exhibiting favorable TPR and FPR, is computationally expensive and unsuitable for high-resolution image data, limiting its applicability.
Table A4: The performance (TPR(SD)/FPR(SD)) on CIFAR-10 and Tiny ImageNet. We mark the best result in boldface while the value with underline denotes the second-best. The failed cases (i.e., TPR < 0.8) are marked in gray. We use the same results as in Table 1 for Adaptive-Blend attack as the 1% and 2% poisoning ratios are sufficiently low. Other attacks have a 5% poisoning ratio.
@@
                                                            CIFAR-10        (All standard deviation values should be multiplied by $10^{-2}$)
Badnet 0.952(6.8)/0.167(0.9) 0.977(1.5)/0.475(0.1) 1.000(0.0)/0.115(0.2) 1.000(0.0)/0.473(0.0) 1.000(0.1)/0.090(1.5) 1.000(0.0)/0.199(0.6)
Blend 0.811(2.3)/0.163(0.0) 0.892(0.5)/0.479(0.0) 0.984(1.5)/0.114(0.3) 1.000(0.1)/0.473(0.0) 0.973(0.5)/0.015(2.6) 0.979(1.1)/0.236(4.4)
TrojanNN 0.951(1.4)/0.175(1.7) 0.925(4.4)/0.478(0.2) 1.000(0.0)/0.117(0.7) 0.969(5.4)/0.475(0.3) 0.998(0.1)/0.018(3.1) 0.967(2.5)/0.201(0.6)
Label-Consistent 1.000(0.0)/0.107(0.8) 0.895(0.6)/0.479(0.0) 0.977(1.0)/0.117(0.8) 0.999(0.1)/0.473(0.0) 0.970(0.8)/0.019(1.7) 0.910(2.0)/0.201(0.4)
WaNet 1.000(0.0)/0.113(0.4) 0.820(0.4)/0.483(0.0) 0.044(1.0)/0.107(0.6) 0.985(1.1)/0.475(0.0) 0.856(0.9)/0.036(2.0) 0.861(6.2)/0.220(4.6)
ISSBA 0.998(0.3)/0.153(4.4) 0.877(1.3)/0.480(0.1) 0.712(7.2)/0.120(0.3) 0.999(0.1)/0.474(0.0) 0.945(1.0)/0.008(1.4) 0.937(3.9)/0.271(5.1)
Adaptive-Blend 0.982(2.6)/0.184(3.6) 0.608(19)/0.145(0.2) 0.014(0.5)/0.069(0.4) 0.753(8.3)/0.144(0.1) 0.000(0.0)/0.023(4.0) 0.779(10)/0.246(5.9)
Average 0.962(1.7)/0.152(1.7) 0.861(3.6)/0.431(0.1) 0.714(1.5)/0.109(0.4) 0.963(1.9)/0.427(0.1) 0.839(0.5)/0.018(2.3) 0.918(3.5)/0.225(3.1)
                                                            Tiny ImageNet    (All standard deviation values should be multiplied by $10^{-2}$)
Badnet 0.996(0.2)/0.093(4.3) 0.452(0.0)/0.502(0.0) 0.878(4.7)/0.109(0.7) 0.548(0.0)/0.497(0.0) / 0.998(0.0)/0.279(1.2)
Blend 0.868(4.2)/0.076(3.8) 0.453(0.0)/0.502(0.0) 0.558(7.9)/0.097(0.4) 0.548(0.0)/0.497(0.0) / 0.624(3.8)/0.269(0.8)
TrojanNN 0.939(2.0)/0.203(0.3) 0.453(0.0)/0.502(0.0) 0.980(2.6)/0.107(0.7) 0.547(0.0)/0.497(0.0) / 0.991(1.2)/0.279(4.3)
Label-Consistent 0.894(15)/0.101(6.9) 0.452(0.0)/0.502(0.0) 0.518(9.1)/0.090(1.2) 0.548(0.0)/0.497(0.0) / 0.978(2.9)/0.092(9.4)
WaNet 0.944(5.2)/0.109(1.2) 0.452(0.0)/0.502(0.0) 0.107(5.1)/0.093(1.1) 0.548(0.0)/0.497(0.0) / 0.314(2.7)/0.267(0.6)
ISSBA 0.794(9.1)/0.217(0.1) 0.452(0.0)/0.502(0.0) 0.994(0.3)/0.104(1.1) 0.547(0.0)/0.497(0.0) / 0.663(1.7)/0.320(0.4)
Adaptive-Blend 0.949(2.9)/0.095(3.7) 0.392(0.1)/0.502(0.0) 0.210(16)/0.099(0.6) 0.621(0.0)/0.497(0.0) / 0.505(13)/0.218(1.2)
Average 0.931(3.7)/0.127(3.3) 0.445(0.0)/0.502(0.0) 0.595(6.8)/0.100(0.8) 0.557(0.0)/0.497(0.0) / 0.684(3.2)/0.265(1.2)
Appendix E Limitations
While this study introduces a promising approach to enhancing the security of DNNs through the Prediction Shift Backdoor Detection (PSBD) method, it primarily focuses on image classification tasks. Image classification represents only a subset of the diverse applications where DNNs are employed, and different data modalities may exhibit unique characteristics and vulnerabilities. Future studies could expand on the PSBD method to encompass a broader range of application domains, potentially adapting the underlying techniques to suit different kinds of neural network architectures and data inputs.
Appendix F Broader Impact
The development of the PSBD method for detecting backdoor attacks in DNNs holds significant potential for positive societal impacts. As DNNs continue to permeate various sectors, including healthcare, autonomous driving, and public safety, ensuring their security is paramount. By improving the reliability of DNNs used in critical infrastructure and safety-critical applications, PSBD helps to prevent potential catastrophic failures that could result from malicious manipulations. In addition, by addressing vulnerabilities in DNNs, this research can help build public trust in AI technologies, encouraging their adoption across more areas of daily life and industry.

### v2

@@ Appendix A Pilot Studies on More Attacks
In this section, we will show the more results of our pilot study in Section 4.1 and a combined method between MC-Dropout[9] and SCP[14]. We maintain the same experimental setup with pilot study that conduct experiments on the CIFAR-10 dataset[23] with ResNet-18[17] trained 100 epochs. Apart from Adaptive-Blend[35] attack, which has a poisoning ratio of 1%, all other attacks maintain a poisoning ratio of 10%.
@@
(a) Benign Model
(b) Blend
(c) TrojanNN
(d) Label-Consistence
(e) ISSBA
(f) Adaptive-Blend
Figure A1: The average MC-Dropout uncertainty of clean training data, backdoor training data and clean validation data under benign and various poisoned models.
@@
In the Section  4.1 of our main paper, we can observe that the model’s mapping from trigger to target label in backdoor data is more salient and robust compared to general image features. In light of this observation, we hypothesize the potential factor of the proximity between backdoor training data uncertainty and clean training data under WaNet[32] attack can be attributed to the absence of sufficient sabotage on the image features. We expect that the model’s ability to extract feature patterns from clean data will be significantly affected, while that from backdoor data will be minimally impacted. Consequently, we can enlarge the uncertainty gap between clean data and backdoor data. One direct method is to increase the dropout rate $p$. However, due to our inability to ascertain the presence or quantity of backdoor data within the suspicious training dataset, we encounter difficulty in determining an optimal value for $p$. Thus, we incorporate elements of SCP approach to introduce a more controllable uncertainty with MC-Dropout uncertainty. SCP can be considered as an instance of input-level uncertainty. It amplifies the pixel values of input images by multiple times, aiming to disrupt the feature patterns in the image. Therefore, we conduct further experiments to see if adding input-level uncertainty further disrupts the general feature patterns in the image.
@@ Results.
(a) Benign Model
(b) BadNets
(c) WaNet
However, under WaNet scenario, one can see from Figure 2(f) that although the uncertainty of the clean training data becomes closer to that of the clean validation data after scaling, the uncertainty gap between clean training and backdoor training data further reduces. Furthermore, as shown in Figure 2(e) and Figure 2(h), under the Label-Consistence[44] and the Adaptive-Blend attack scenarios, the backdoor training data exhibits only slightly smaller uncertainty than that of clean training data, which poses significant challenges for their differentiation.
Appendix B Prediction Shift Phenomenon on More Scenarios
B.1 PS Phenomenon on More Poisoned Models.
Except for the ISSBA[25] attack, the shift intensity of clean data is not pronounced. In other attacks, the shift intensity of clean data is extremely strong. More importantly, in all attack scenarios, clean data exhibits a bias towards the target class (class 0 in our experiments). This further indicates that when the model has good generalizability, the neuron bias path established by the backdoor data in the model becomes more stable and specific. This property may be exploited in the future to detect the target class of backdoor attacks.
B.2 PS Phenomenon on Tiny ImageNet Dataset.
From the observation of Figure A4, we can find that the PS phenomenon and neuron bias effect persist even in the more complex Tiny ImageNet dataset[37].
@@ Appendix B Prediction Shift Phenomenon on More Scenarios
Despite assuming that the defender can freely choose the model architecture, we also conducted experiments using the VGG16-bn[39] model to show that our method is not dependent on any specific model architecture. The experiment was conducted as follows Section 4.2.
@@
Appendix C Detailed Results of Neuron Bias Effect
In this section, we present the more complete results of the “neuron bias” effect on the BadNets and WaNet models. As illustrated in Figures A6 and A7, there is a pronounced and widespread presence of the neuron bias effect. This confirms our hypothesis that the PS Phenomenon is a result of the neuron bias effect.
(a) Clean Image
(b) Backdoor Image
Figure A6: The all 512 activation maps extracted by the top layer of the BadNets model with dropout. The red boxes represent the feature map values are non-zero and the difference between each activation value in the clean and backdoor feature maps is no greater than 1. The features of clean and backdoor image become almost identical with dropout, verifying the existence of neuron bias effect.
(a) Clean Image
(b) Backdoor Image
Figure A7: The all 512 activation maps extracted by the top layer of the WaNet model with dropout. The features of clean and backdoor image become more identical with dropout, verifying the neuron bias effect is not limited to specific attack.
Appendix D Experiments Details
D.1 Detailed Settings for Datasets and Training of Backdoored Models.
@@ Appendix D Experiments Details
Dataset # Input size Classes Training images Testing images
CIFAR-10 3$\times$32$\times$32 10 50,000 10,000
GTSRB 3$\times$32$\times$32 43 39,209 12,630
Tiny ImageNet 3$\times$64$\times$64 200 100,000 10,000
@@
CA ASR   CA ASR   CA ASR   CA ASR   CA ASR   CA ASR   CA ASR   CA
GTSRB 0.981 1.000   0.982 0.999   0.975 1.000   0.980 0.994   0.975 0.988   0.982 0.999   0.983 0.913   0.952
The details of datasets and training procedures of DNN models in our experiments are summarized in Table A1 and Table A2. Note that clean validation data refers to the clean data we used to filter backdoor training data, which was 5% of the total quantity of whole training dataset randomly selected from the test set of CIFAR-10 and GTSRB[40], and the validation set of Tiny ImageNet, respectively.
D.2 Implements of Backdoor Attacks.
BadNets, TrojanNN[28], and Blend[6] correspond to typical all-to-one label-poisoned attacks with patch-like trigger, generated trigger, and blending trigger respectively. Label-Consistent is a representative clean label attack. WaNet is an image transformation-based invisible attack. ISSBA is an effective sample-specific invisible attack which generates sample-specific invisible additive noises as backdoor triggers. It generates sample-specific invisible additive noises as backdoor triggers by encoding an attacker-specified string into benign images through an encoder-decoder network. Adaptive-Blend is an adaptive poisoning strategy suggested that can suppress the latent separability characteristic.
@@
We implemented this attack using the backdoor-toolbox. The trigger we used on CIFAR-10 and GTSRB is a 3$\times$3 checkerboard placed in the bottom right corner of an image, and a 6$\times$6 trigger placed in the same position on Tiny ImageNet.
@@
On CIFAR-10, we directly use the adversarial images provided by the original paper;  [^3] https://github.com/MadryLab/label-consistent-backdoor-code on GTSRB and Tiny ImageNet we use the adversarial images provided by BackdoorBench.
@@
We implemented this attack using the backdoor-toolbox. Following the original paper [35], we choose “Hello Kitty” trigger and set the the cover ratio equal to the poisoning ratio. Compared to the original paper, we only added the cover ratio and poisoning ratio to ensure that the attack success rate exceeds 85%. On CIFAR-10 and GTSRB, we selected the “Hello Kitty” trigger, setting both the cover ratio and poisoning ratio to 0.01, and the blend ratio to 0.2. On Tiny ImageNet, we chose a random noise trigger, setting both the cover ratio and poisoning ratio to 0.02, and the blend ratio to 0.15.
D.3 Implements of Baseline Defences.
We implement Spectral Signature [43], Strip[11], Spectre[16] and SCAN [42] based on the original implementation provided by the backdoor-toolbox. We have implemented the SCP [14] based on the backdoor-toolbox. We use the original implementation of CD-L [19] and follow the hyperparameter settings specified in the original paper.
D.4 Performance of the Benign and Poisoned Models.
Consistent with the methodology employed in previous backdoor attack studies, we utilize performance metrics to assess the effectiveness of the backdoor attacks: attack success rate (ASR) and clean accuracy (CA). ASR denotes the success rate in classifying the poisoned samples into the corresponding target classes. CA measures the accuracy of the backdoored model on the benign test dataset. ASR and CA for different backdoor attacks are included in Table A3. The poisoning ratio is 10%.
D.5 Computational Environment
@@
Appendix E Detection Performance on Low Poisoning Ratio Scenario
@@
As shown in Table A4, PSBD demonstrated superior performance against various backdoor attacks under low poisoning scenarios, outperforming all other baseline methods on average. However, its performance exhibited a slight decline in certain attack settings. This degradation may be attributed to the reduced robustness of neuron bias paths within the model due to the limited amount of backdoor data, making it more susceptible to random fluctuations. The performance of other baseline methods deteriorated significantly, particularly on the more challenging Tiny ImageNet dataset.
Table A4: The performance (TPR/FPR) on CIFAR-10, GTSRB and Tiny ImageNet. We mark the best result in boldface while the value with underline denotes the second-best. The failed cases (i.e., TPR < 0.8) are marked in gray. We use the same results as in Table 1 for Adaptive-Blend attack as the 1% and 2% poisoning ratios are sufficiently low. Other attacks have a 5% poisoning ratio. OOT indicates that the method did not finish within the allocated time limit.
@@
                 CIFAR-10
Badnet 0.979/0.158 0.977/0.475 1.000/0.115 1.000/0.473 1.000/0.090 1.000/0.199 0.999/0.164
Blend 0.899/0.176 0.892/0.479 0.984/0.114 1.000/0.473 0.973/0.015 0.979/0.236 0.957/0.161
TrojanNN 0.951/0.175 0.925/0.478 1.000/0.117 0.969/0.475 0.998/0.018 0.967/0.201 1.000/0.162
Label-Consistent 1.000/0.107 0.895/0.479 0.977/0.117 0.999/0.473 0.970/0.019 0.910/0.201 0.994/0.166
WaNet 1.000/0.113 0.820/0.483 0.044/0.107 0.985/0.475 0.856/0.036 0.861/0.220 0.430/0.149
ISSBA 0.998/0.153 0.877/0.480 0.712/0.120 0.999/0.474 0.945/0.008 0.937/0.271 0.984/0.163
Adaptive-Blend 0.982/0.184 0.608/0.145 0.014/0.069 0.753/0.144 0.000/0.023 0.779/0.246 0.432/0.167
Average 0.973/0.152 0.861/0.431 0.714/0.109 0.963/0.427 0.839/0.018 0.918/0.225 0.828/0.162
                 GTSRB
Badnet 0.993/0.202 0.448/0.502 0.995/0.094 0.552/0.497 OOT 1.000/0.328 0.839/0.188
Blend 0.859/0.223 0.448/0.502 0.915/0.094 0.552/0.497 OOT 0.301/0.334 0.072/0.198
TrojanNN 0.978/0.206 0.449/0.502 0.408/0.093 0.551/0.497 OOT 0.150/0.329 0.515/0.191
Label-Consistent 0.844/0.202 0.449/0.502 0.998/0.113 0.551/0.497 OOT 0.956/0.403 0.198/0.172
WaNet 0.999/0.085 0.448/0.502 0.030/0.102 0.552/0.497 OOT 0.043/0.320 0.022/0.185
ISSBA 0.986/0.214 0.449/0.502 0.469/0.102 0.551/0.497 OOT 0.590/0.334 0.446/0.195
Adaptive-Blend 0.899/0.194 0.299/0.392 0.004/0.094 0.750/0.388 OOT 0.071/0.332 0.028/0.158
Average 0.937/0.189 0.427/0.486 0.546/0.099 0.580/0.481 OOT 0.444/0.340 0.303/0.184
                 Tiny ImageNet
Badnet 0.996/0.093 0.452/0.502 0.878/0.109 0.548/0.497 OOT 0.998/0.279 0.390/0.178
Blend 0.871/0.065 0.453/0.502 0.558/0.097 0.548/0.497 OOT 0.624/0.269 0.376/0.185
TrojanNN 0.939/0.203 0.453/0.502 0.980/0.107 0.547/0.497 OOT 0.991/0.279 0.990/0.166
Label-Consistent 0.983/0.100 0.452/0.502 0.518/0.090 0.548/0.497 OOT 0.978/0.092 0.967/0.154
WaNet 0.944/0.109 0.452/0.502 0.107/0.093 0.548/0.497 OOT 0.314/0.267 0.403/0.150
ISSBA 0.890/0.216 0.452/0.502 0.994/0.104 0.547/0.497 OOT 0.663/0.320 0.644/0.140
Adaptive-Blend 0.949/0.095 0.392/0.502 0.210/0.099 0.621/0.497 OOT 0.505/0.218 0.331/0.176
Average 0.939/0.126 0.445/0.502 0.595/0.100 0.557/0.497 OOT 0.684/0.265 0.586/0.164
Appendix F Resistance to Potential Adaptive Attacks
The most common adaptive attack scenario is one with a low poisoning ratio. As shown in Section E, our PSBD method demonstrates effective performance.
We further evaluate the robustness of PSBD against potential adaptive attacks in the worst-case scenario, where adversaries have complete knowledge of our defense. Typically, a vanilla backdoored model performs normally with benign samples but produces adversary-specific predictions when exposed to poisoned samples. The objective function for training such a model with a poisoned training dataset can be represented as follows:
$\min_{\boldsymbol{\theta}}\mathcal{L}_{bd}(\mathcal{D}^{c}\cup\mathcal{D}^{b};\boldsymbol{\theta})$ (3)
where $\boldsymbol{\theta}$ denotes the model parameters and $\mathcal{L}$ is the cross entropy loss function. We develop an adaptive attack by introducing a loss term specifically designed to ensure that benign samples have a low PSU value. This adaptive loss item $\mathcal{L}_{ada}$ is defined as:
$\mathcal{L}_{ada}=\ \phi_{PSU}(\mathbf{x}),\mathbf{x}\in\mathcal{D}^{c}\cup\mathcal{D}^{b}$ (4)
Subsequently, we integrate this adaptive loss $\mathcal{L}_{ada}$ with the vanilla loss $\mathcal{L}_{bd}$ to formulate the overall loss function as $\mathcal{L}=(1-\alpha)\mathcal{L}_{bd}+\alpha\mathcal{L}_{ada}$ , where $\alpha$ is a weighting factor. We then optimize the original model’s parameters $\boldsymbol{\theta}$ by minimizing $\mathcal{L}$ during the training phase. Please note that, to maintain the effectiveness of the training process, we compute $\mathcal{L}_{ada}$ every 50 iterations. As in previous experiments, we also use two representative backdoor attacks, BadNets and WaNet, to develop adaptive attacks on the CIFAR-10 dataset.
The adversary aims to find a value of $\alpha$ that best balances the ASR and CA. Table A5 presents the performance of
Table A5: The attack performance of adaptive attacks.
$\alpha$rightarrow 0.2 0.5 0.9
[…]
Badnet 1.000 0.828 1.000 0.838 1.000 0.836
WaNet 0.899 0.803 0.931 0.820 0.925 0.823
Table A6: Performance (TPR/FPR) of PSBD under adaptive attacks.
$\alpha$rightarrow 0.2 0.5 0.9
[…]
Badnet 0.989/0.157 0.997/0.131 0.967/0.127
WaNet 1.000/0.119 0.998/0.138 1.000/0.114
the adaptive attacks under various $\alpha$ settings. As shown in the results, both attacks (BadNets and WaNet) on the CIFAR-10 dataset consistently exhibit high ASR and CA across different values of $\alpha$, highlighting the effectiveness of the adaptive attacks.
On the other hand, Table A6 shows that adaptive attacks can still be effectively defended by our method. We conducted further investigation and observed that the adaptive loss indeed caused the model to behave differently from the non-adaptive version. However, our defense can adapt to modified backdoor models. Specifically, we observed that a dropout rate of 0.7 was used for the non-adaptive backdoor model, as detailed in Section 4.2 of the main paper. In contrast, the dropout rate for the adaptive model was 0.8. In other words, our algorithm learns to select the appropriate dropout rate for different models. This ability to counter adaptive attacks is a key advantage of our method compared to previous approaches.
Table A7: The AUROC values (AUROC) on Tiny ImageNet.
[…]
Badnet 0.993 0.474
Blend 0.958 0.922
TrojanNN 0.963 0.877
Label-Consistent 0.996 0.548
WaNet 0.985 0.831
ISSBA 0.907 0.801
Adaptive-Blend 0.969 0.705
Average 0.967 0.737
Appendix G AUROC Metric
In addition to the TPR and FPR metrics, we also compare the AUROC metric with the CD-L method. We evaluate the AUROC on the more challenging Tiny ImageNet dataset, and the results are presented in Table A7. We observe that our method achieves a high AUROC score across various attacks, which verifies the robustness of our method in selecting the threshold parameter $T$.
Appendix H Limitations
While this study introduces a promising approach to enhancing the security of DNNs through the PSBD method, it also has several limitations.
The majority of existing approaches, such as SCP, CatchBackdoor [21] and our PSBD primarily rely on empirical findings with limited theoretical foundations. Developing solid theoretical justifications for these methods remains important future work.
Our experiments use model architectures and datasets consistent with prior studies [14, 21] to ensure fair comparability, which represents the most common experimental setup in the field. As discussed in Sections 3.2 and 5.1, while defenders can employ any model or training strategy to effectively detect backdoor data, the generalizability of the Prediction Shift phenomenon to more complex architectures, such as ViT[8], and larger-scale datasets remains a valuable avenue for future exploration.